Please help me understand this problem. A baseball team's manager buys 70 balls for $260. The balls used for practice cost $3 each and the balls used in games cost $4 each. How many balls used in games did the manger buy?
Answer by rfer(16322)

3p+4g=260
p+g=70
p=70-g
------------
3(70-g)+4g=260
210-3g+4g=260
g=50 game balls
----------------
p=70-g
p=20
----------
3*20+4*50=260
60+200=260